# Assembly Theory is an approximation to algorithmic complexity based on LZ compression that does not explain selection or evolution

### v5

@@ Abstract
We demonstrate that Assembly Theory (AT) is reliant on Shannon Entropy and a method based upon statistical compression that belongs to the LZ family of compression algorithms, family which comprises popular algorithms used in ZIP, GZIP, or JPEG. Such popular algorithms have been shown to empirically reproduce the results of AT that were reported before in successful applications to separating organic from non-organic molecules and in the context of the study of selection and evolution. In particular, the assembly index calculation method is an LZ compression scheme, even if used for classification or explanatory purposes. We show that the assembly index value is equivalent to the size of a minimal context-free grammar. The statistical compressibility of such a method is bounded by Shannon Entropy and other equivalent traditional LZ compression schemes, such as LZ77, LZ78, or LZW. In addition, we demonstrate that AT, and the algorithms supporting its pathway complexity, assembly index, and assembly number, define compression schemes and methods that are subsumed into the theory of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity. Due to AT’s current lack of logical consistency in defining causality for non-stochastic processes and lack of the empirical evidence that it outperforms other complexity measures found in the literature capable of explaining the same phenomena, we conclude that the assembly index and the assembly number do not lead to an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by (abiotic or Darwinian) selection and evolution that could not have been made with Shannon Entropy or had not been reported before with algorithmic complexity.
@@
assembly theory, assembly index/number, pathway complexity, algorithmic complexity, lossless compression, LZ compression, LZ77, LZ78, logical depth, selection, evolution.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of AT advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. This has been contested in [3], whose results “demonstrate that abiotic chemical processes have the potential to form crystal structures of great complexity”, exceeding the assembly index (or MA, when applied to quantify the assembly index on molecules) threshold proposed by AT’s authors. The existence of such abiotic objects would lead AT’s methods to be prone to false positives, corroborating the predictions in [4]. The assembly index (or MA) was shown to perform equally, if not worse in some cases, to popular compression algorithms (including those of statistical nature) [4], some of which have been applied before in [5, 6, 7].
The forthcoming sections show the connections of AT to well-established theories that AT’s authors claim are different from and unrelated to [8]. We show that the claims advanced by its authors that AT unifies life and biology with physics [2, 9] rely on a circular argument and the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution not already known (hierarchical modularity), already made (in connection to complexity) [6, 10, 11], or previously investigated by information- and graph-theoretic approaches to chemical and molecular complexity [12, 13, 14, 5].
@@
In addition to the assembly index calculation method being a CFG-based compression method bounded by traditional LZ schemes, we demonstrate that the assembling process that results in the construction of an object is a LZ scheme. The compressibility achieved by this LZ scheme depends on the length of the shortest assembling paths and the “simplicity” of the minimal assembly spaces. This notion of simplicity encompasses both how the assembly space structure differs from a single-thread (or linear) space and how many more distinct assembling paths can lead to the same object. Such a LZ scheme reduces the complexity (which the assembly index aims to quantify) of the assembling process of an object to compressibility and computational resource efficiency of a compression method.
@@
The theory and methods of AT describing a compression algorithm are, therefore, subsumed into Algorithmic Information Theory (AIT), which in turn has been applied in the same areas that AT has covered before, from organic versus non-organic compound classification to bio and technosignature detection and selection and evolution [5, 16, 6, 17]. This makes the assembly index and assembly number purported by AT effectively approximations to algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity [18, 19, 20, 21].
@@
As proposed in [2], the assembly number is intended to measure the amount of selection and evolution necessary to produce the ensemble of (assembled) objects. That is, the assembly number aims to quantify the presence of constraints or biases in the underlying generative processes (e.g., those part of the environment in which the objects were assembled) of the ensemble, processes that set the conditions for the appearance of the assembled objects. A higher assembly number—not to be conflated with the assembly index—would mean that more “selective forces” were put into play as e.g. environmental constraints or biases in order to allow or generate a higher concentration of high-assembly-index elements. Otherwise, these high-assembly-index objects would not occur as often in the ensemble.
@@
Consonant with the constraints and biases that the assembly number aims to quantify but that, in fact, we demonstrate it constitutes a compression method subsumed into AIT, ensembles with higher assembly numbers are more compressible and, therefore, would diverge more from those ensembles that are more statistically random with a lesser presence of constraints and biases. If one assumes that the assembly number quantifies the presence of constraints or biases in the underlying generative processes of the ensembles, then a more compressible ensemble implies that more constraints or biases played a role in generating more high-assembly-index objects more frequently than would have been the case in an environment with otherwise fewer constraints and biases (i.e., a more random or incompressible environment), hence increasing the frequency of occurrence of high-assembly-index (i.e., less compressible) objects in this environment.
@@
Conversely, under the same assumption, in case more biotic processes are present in an ensemble, it implies a higher assembly number, and thereby it would also imply that the ensemble is more compressible. For example, this occurs in scenarios where there is a stronger presence of top-down (or downward) causation [11] behind the possibilities or paths that lead to the construction of the objects, while a less compressible ensemble would indicate a weaker presence (or absence) of top-down causation.
@@
We, therefore, conclude that AT cannot offer a different or better explanation of selection, evolution, or top-down causation than the connections already established [6, 11] consistent with our previous position that a single scalar is unlikely to classify life or quantify selection or evolution independent of the environment and the perturbations such environment imposes on the objects and the assembly process.
@@ 2.1 The assembly index is a compression algorithm of the LZ family
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [25, 26]. These more advanced measures have found application in the same contexts and domains explored by AT, encompassing tasks like distinguishing organic from non-organic molecules [5], investigating potential connections to selection and evolution [6, 10], the detection of bio and techno signatures [27, 28], and explorations into causality [29, 30]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Sup. Inf. A and [4]). Thus, it becomes evident that AT and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
@@
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Suplementary Information and [4]. This alignment with ‘Template Program A’ effectively highlights the association of AT with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to present a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) [31] further underscores the disconnection of AT and their authors’ suggestion to reinvent text compression (e.g. ZIP or GZIP) from the current state of the art in the field of statistical and non-statistical compression beyond LZW.
It is worth noting that although there may be minor variations in the implementation details of the assembly index and the assembly number of which the authors themselves have proposed significantly different versions, we demonstrate that they are all qualitatively and quantitatively equivalent to the LZ compression algorithms introduced in the 1970s, such as LZ77/LZ78. See Sup. Inf. A.
In essence, the assembly index is fundamentally underpinned by the LZ encoding of the objects it measures. This reveals that the assembly index, and consequently AT, aligns more closely with the principles of traditional information theory (Shannon Entropy), and statistical data compression than the authors are willing to acknowledge [8]. In fact, both AT and the statistical compression methods that underpins it are subsumed into the theory of algorithmic complexity.
The AT authors also assert that their index’s ability to differentiate between organic and non-organic compounds validates its natural applicability. However, when compared with other statistical indexes, including various compression algorithms, these alternative methods often engender similar or superior performances [4]. This undermines the claims made in favour of AT.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The argument presented by the proponents of AT regarding the uncomputability of algorithmic complexity is deeply misguided. While it is true that $\mathbf{K}$ is semi-computable, computable algorithms like LZ77/LZ78/LZW have been widely used to approximate it. These algorithms have been applied in biology and chemistry, challenging the assertion that AT represents a unique or superior approach  [5, 4]. Furthermore, in an earlier 1976 article [35], Lempel and Ziv defined an early version of their algorithm directly as a computable method to approximate algorithmic complexity, defining what is known as LZ complexity, which speaks to the initial motivation behind their now seminal compression algorithm.
@@
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is a misdirected concern. A Turing machine is a synonym of ‘algorithm’ and an ‘algorithm’ is a synonym of a Turing machine, any rule or method that is algorithmic in nature, including all the computable methods in AT, are algorithms, and, therefore, technically, Turing machines or programs that can run on a (universal) Turing machine. These facts are only a technicality unrelated to any putative advantage of AT over other approaches to algorithmic complexity.
@@
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are, in fact, established because of the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT [31] indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Sup. Inf. section A made any mention of any Turing machine. In the same manner, all the mentions to Turing machines in [4] can also be equivalently replaced by any sufficiently expressive programming language running on an arbitrary computer.
@@
In practice, computable approximations like LZ77/LZ78—which AT mimics without attribution—do not require a Turing machine, and have been widely and successfully used in clustering and categorisation  [32, 20], including in the successful separation of chemical compounds into organic and non-organic categories [5]. To say that a physical (computable/recursive) process like those affected by AT’s methods is not a compression algorithm because it does not resemble or correspond to the functioning of a Turing machine is as naively wrong and misplaced as saying a program written in Python cannot model the movement of a pendulum because nature does not run Python.
@@
This type of argument also reflects a misunderstanding of the purposes of Turing machines and the foundations of computer science. As explained in a quote often attributed to Edsger W. Dijkstra,
@@
In a published paper [8], the authors offered a proof of computability of their assembly index to distance themselves from Shannon Entropy and algorithmic (Kolmogorov) complexity. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability but also all other resource-bounded approximations to $\mathbf{K}$ are computable, including LZW that has been used for decades for this purpose [20].
2.3 Conflating object assembly process and causation directionality
The authors assume and present the sequential nature of their algorithm as an advantage [31]. Assembly Theory claims to be able to extract causal knowledge in the form of chains of cause and effect from its assembly pathways [31]. The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled. This is no different from, and is indeed a restricted version of, a pushdown automaton capable of instantiating grammar compressor (which the authors rename as ‘assembly pathway’) and works exactly like a LZ algorithm. See Sup. Inf. A.
@@ 2.3 Conflating object assembly process and causation directionality
This assumption of sequential assembly of an object based on causal direction is manifestly incorrect regarding how an object subject to selection assembles. In contrast, only by finding the generative mechanism of the object—such as the underlying set of (computable) mechanisms (of which the assembly index only takes into account one instance, the program that only counts copies)—and thus explaining the object in a non-trivial fashion, can one reproduce both the causal direction from the sequence of connected steps (see Fig. 1B) as well as how the object itself may have been assembled (which is not and cannot be in a sequential fashion).
@@
The authors of AT have mistakenly claimed that algorithmic complexity and Turing machines are not, or cannot be, related to causality. This is incorrect [29, 36]. They were introduced as causal artifacts for studying mechanistic means for logical operations.
@@
The assembly index calculation from the minimum rooted assembly space being reduced to a grammar compression scheme demonstrates not only that the assembly process’ complexity is dependent on the contingencies regarding which assembly spaces, or which distinct paths, were taken in the historical past stages to build the object, but also that this recursivity and dependence on past trajectories are subsumed into the concept of compressibility. However, unlike algorithmic approaches, assembly theory does not incorporate any of the elements of causality such as perturbation and counterfactual analysis, as has been done in the context of algorithmic complexity [30], a generalisation that compared to AT reflects the simplistic nature of AT.

### v6

@@ Abstract
We prove the full equivalence between Assembly Theory (AT) and Shannon Entropy via a method based upon the principles of statistical compression renamed ‘assembly index’ that belongs to the LZ family of popular compression algorithms (ZIP, GZIP, JPEG). Such popular algorithms have been shown to empirically reproduce the results of AT, results that have also been reported before in successful applications to separating organic from non-organic molecules and in the context of the study of selection and evolution. We show that the assembly index value is equivalent to the size of a minimal context-free grammar. The statistical compressibility of such a method is bounded by Shannon Entropy and other equivalent traditional LZ compression schemes, such as LZ77, LZ78, or LZW. In addition, we demonstrate that AT, and the algorithms supporting its pathway complexity, assembly index, and assembly number, define compression schemes and methods that are subsumed into the theory of algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity. Due to AT’s current lack of logical consistency in defining causality for non-stochastic processes and the lack of empirical evidence that it outperforms other complexity measures found in the literature capable of explaining the same phenomena, we conclude that the assembly index and the assembly number do not lead to an explanation or quantification of biases in generative (physical or biological) processes, including those brought about by (abiotic or Darwinian) selection and evolution, that could not have been arrived at using Shannon Entropy or that have not been reported before using classical information theory or algorithmic complexity.
@@
assembly theory, assembly index/number, pathway complexity, algorithmic complexity, lossless compression, LZ compression, LZ77, LZ78, logical depth, selection, evolution, Block Decomposition Method, algorithmic probability.
@@ 1 Introduction
Assembly Theory (AT) has recently garnered significant attention in some academic circles and the scientific media. Responding to how open-ended forms can emerge from matter without a blueprint, AT purports to explain and quantify selection and evolution. The central claim of AT advanced in [1, 2] is that objects with a high assembly index “are very unlikely to form abiotically”. This has been contested in [3], whose results “demonstrate that abiotic chemical processes have the potential to form crystal structures of great complexity”, exceeding the assembly index threshold (or MA, when applied to quantify the assembly index on molecules) proposed by AT’s authors. The existence of such abiotic objects would render AT’s methods prone to false positives, corroborating the predictions in [4]. The assembly index (or MA) was shown to perform equally well, or worse in some cases, relative to popular compression algorithms (including those of a statistical nature) [4], some of which have been applied before in [5, 6, 7].
The following sections show the connections of AT to well-established theories that AT’s authors claim are different from and unrelated to [8]. We show that the claim advanced by its authors that AT unifies life and biology with physics [2, 9] relies on a circular argument and on the use of a popular compression algorithm, with no empirical or logical support to establish deeper connections to selection and evolution than those already known (hierarchical modularity), already made (in connection to complexity) [6, 10, 11], or previously investigated using information- and graph-theoretic approaches to chemical and molecular complexity [12, 13, 14, 5].
@@
In addition to the assembly index calculation method being a CFG-based compression method bounded by traditional LZ schemes, we demonstrate that the assembling process that results in the construction of an object is a LZ scheme. The compressibility achieved by this LZ scheme depends on the length of the shortest assembling paths and the “simplicity” of the minimal assembly spaces. This notion of simplicity encompasses both how the assembly space structure differs from a single-thread (or linear) space and how many more distinct assembling paths can lead to the same object. Such a LZ scheme reduces the complexity (which the assembly index aims to quantify) of the assembling process of an object to the compressibility and computational resource efficiency of a compression method.
@@
The theory and methods of AT describing a compression algorithm are, therefore, subsumed into Algorithmic Information Theory (AIT), which in turn has been applied in the same areas that AT has covered, from organic versus non-organic compound classification to bio- and technosignature detection and selection and evolution [5, 16, 6, 17]. This makes the assembly index and assembly number proposed by AT effectively approximations to algorithmic (Kolmogorov-Solomonoff-Chaitin) complexity [18, 19, 20, 21].
@@
As proposed in [2], the assembly number is intended to measure the amount of selection and evolution necessary to produce the ensemble of (assembled) objects. That is, the assembly number aims to quantify the presence of constraints or biases in the underlying generative processes (e.g., those parts of the environment in which the objects were assembled) of the ensemble, processes that set the conditions for the appearance of the assembled objects. A higher assembly number—not to be conflated with the assembly index—would mean that more “selective forces” were in play as, e.g., environmental constraints or biases, in order to allow or generate a higher concentration of high-assembly-index elements. Otherwise, these high-assembly-index objects would not occur as often in the ensemble.
@@
Consonant with the constraints and biases that the assembly number aims to quantify, though in fact, as we demonstrate, it constitutes a compression method subsumed into AIT, ensembles with higher assembly numbers are more compressible and would therefore diverge more markedly from those ensembles that are more statistically random, having fewer constraints and biases. If one assumes that the assembly number quantifies the presence of constraints or biases in the underlying generative processes of the ensembles, then a more compressible ensemble implies that more constraints or biases played a role in generating more high-assembly-index objects more frequently than would have been the case in an environment with fewer constraints and biases (i.e., a more random or incompressible environment), thus increasing the frequency of occurrence of high-assembly-index (i.e., less compressible) objects in this environment.
@@
Conversely, under the same assumption, the presence of more biotic processes in an ensemble implies a higher assembly number, which in turn would imply that the ensemble is more compressible. This occurs, for example, in scenarios where there is a stronger presence of top-down (or downward) causation [11] behind the possibilities or paths that lead to the construction of the objects, while a less compressible ensemble would indicate a weaker presence (or absence) of top-down causation.
@@
We therefore conclude that AT cannot offer a different or better explanation of selection, evolution, or top-down causation than the connections already established [6, 11], consistent with our previous position that a single scalar is unlikely to classify life or quantify selection or evolution independent of the environment and the perturbations it imposes on the objects and on the assembly process.
@@ 2.1 The assembly index is a compression algorithm of the LZ family
In contrast, more robust approximations to $\mathbf{K}$, capable of capturing blocks and other causal content within an object, have been proposed for purposes ranging from exploring cause-and-effect chains to quantifying object memory and characterising process content [25, 26]. These more advanced measures have found application in the same contexts and domains explored by AT, encompassing tasks like distinguishing organic from non-organic molecules [5], investigating potential connections to selection and evolution [6, 10], the detection of bio- and technosignatures [27, 28], and explorations into causality [29, 30]. Importantly, when applied to the data employed as evidence by AT, the more sophisticated measures consistently outperform the assembly index (see Sup. Inf. A and [4]). Thus, it becomes evident that AT and its assembly index represent a considerably constrained version of compression algorithms, and a loose upper bound of $\mathbf{K}$.
@@
Any approximation to $\mathbf{K}$ that accounts for identical repetitions, including all known lossless statistical compression algorithms, can achieve equivalent or superior results, as demonstrated in the Supplementary Information and [4]. This alignment with ‘Template Program A’ effectively highlights the association of AT with well-established principles of compression and coding theory, thereby refuting the initial claim of its authors to present a unique methodology. Moreover, the authors’ suggestion that their index may be generalised as a universally applicable algorithm for any object (including text) [31] further underscores the disconnect between AT– and its authors’ drive to reinvent traditional algorithms such as text compression based on Shannon Entropy–and the current state of the art in the field of statistical and non-statistical compression beyond LZW [30].
It is worth noting that although there may be minor variations in the implementation details of the assembly index and the assembly number, of which the authors themselves have proposed significantly different versions, we demonstrate that they are all qualitatively and quantitatively equivalent to the LZ compression algorithms introduced in the 1970s, such as LZ77/LZ78. See Sup. Inf. A.
In essence, the assembly index is fundamentally underpinned by the LZ encoding of the objects it measures. This reveals that the assembly index, and consequently AT, aligns more closely with the principles of traditional information theory (Shannon Entropy), and statistical data compression than the authors are willing to acknowledge [8]. In fact, both AT and the statistical compression methods that underpin it are subsumed into the theory of algorithmic complexity.
The AT authors also assert that their index’s ability to differentiate between organic and non-organic compounds validates its natural applicability. However, when compared with other statistical indexes, including various compression algorithms, these alternative methods often result in similar or superior performances [4]. This undermines the claims made in favour of AT.
@@ 2.2 A compression algorithm is a mechanical procedure that corresponds to a physical process
The argument presented by the proponents of AT regarding the uncomputability of algorithmic complexity is deeply misguided. While it is true that $\mathbf{K}$ is semi-computable, computable algorithms like LZ77/LZ78/LZW have been widely used to approximate it. These algorithms have been applied in biology and chemistry, challenging the assertion that AT represents a unique or superior approach  [5, 4]. Furthermore, in a 1976 article [35], Lempel and Ziv defined an early version of their algorithm directly as a computable method to approximate algorithmic complexity, defining what is known as LZ complexity, which speaks to the initial motivation behind their now seminal compression algorithm.
@@
The emphasis on the purported requirement of a Turing machine in the context of algorithmic complexity is a misdirected concern. ’Turing machine’ is synonymous with ‘algorithm’ and an ’algorithm’ is a synonym for a ’Turing machine’. Any rule or method that is algorithmic in nature, including all the computable methods in AT, are algorithms, and are therefore technically Turing machines or programs that can run on a (universal) Turing machine. These facts are only a technicality unrelated to any putative advantage of AT over other approaches to algorithmic complexity.
@@
On the contrary, one of the distinctive features of algorithmic information theory (AIT), algorithmic complexity being one of its indexes, is that its importance and pervasiveness in mathematics, theoretical computer science, and complexity science are in fact owed to the invariance of its results with respect to the model of computation, whether abstract or physically implemented. More specifically, such statements from the authors of AT [31] indicate a fundamental misunderstanding of equivalence classes closed under reductions (in this case, computability classes), which are basic and pivotal concepts in computer science, complexity theory, and mathematics in general. It is evidence in support of this point that none of the formal proofs in the Sup. Inf. section A make any mention of any Turing machine. In the same manner, all Turing machines referenced in [4] can be replaced by any sufficiently expressive programming language running on an arbitrary computer.
@@
In practice, computable approximations like LZ77/LZ78—which AT mimics without attribution—do not require a Turing machine, and have been widely and successfully used in clustering and categorisation  [32, 20], including in the successful separation of chemical compounds into organic and non-organic categories [5], and in the reconstruction of evolutionary phylogenetic trees [32]. To say that a physical (computable/recursive) process like those affected by AT’s methods is not a compression algorithm because it does not resemble or correspond to the functioning of a Turing machine is as naively wrong and misplaced as saying a program written in Python cannot model the movement of a pendulum because nature does not run Python. A Turing machine is an abstraction and a synonym of an ‘algorithm.’ Everything is an algorithm in Assembly Theory, and therefore it is governed by the same principles of computer science and information theory.
@@
This type of argument also reflects a misunderstanding of the purposes of Turing machines and of the foundations of computer science. As explained in a quote often attributed to Edsger W. Dijkstra,
@@
In a published paper [8], the authors offered a proof of computability of their assembly index to distance themselves from Shannon Entropy and algorithmic (Kolmogorov) complexity. Their algorithm, categorised as Template Program A, is trivially computable and requires no proof of computability, but all other resource-bounded approximations to $\mathbf{K}$ are also computable, including LZW that has been used for 60 years for similar purposes [20].
2.3 Conflating object assembly process and directionality of causation
The authors assume and present the sequential nature of their algorithm as an advantage [31]. Assembly Theory (AT) claims to be able to extract causal knowledge by measuring the degree of causality in the form of non-unique chains of cause and effect across the assembly pathways [31]. To tackle the problem of top-down (or downward) causation [11], its assembly number should be able to “detect the emergence of new levels of organisation and their causal influence on lower-level phenomena in the world you are observing” [36].
The authors’ central assumption is that each step of the basic Template Program A constitutes a cause-and-effect chain corresponding to how an object may have been physically assembled from an assembly path, which may not be unique. This is no different from, and is indeed a restricted version of, a (deterministic or non-deterministic) pushdown automaton capable of instantiating a grammar compressor (which the authors rename an ‘assembly pathway’) and works exactly like a LZ algorithm. See Sup. Inf. A.
@@ 2.3 Conflating object assembly process and directionality of causation
This assumption of sequential assembly of an object based on causal direction is manifestly incorrect in regards to how an object subject to selection assembles. In contrast, only by finding the generative mechanism of the object—such as the underlying set of (computable) mechanisms (of which the assembly index only takes into account one instance, the program that only counts copies)—and thus explaining the object in a non-trivial fashion, can one reproduce both the causal direction from the sequence of connected steps (see Fig. 1B) and how the object itself may have been assembled (which is not and cannot be in a sequential fashion).
@@
The authors of AT have mistakenly claimed that algorithmic complexity and Turing machines are not, or cannot be, related to causality. This is incorrect [29, 37]. They were introduced as causal artifacts for studying mechanistic means of performing logical operations.
@@
That the calculation of the assembly index from the minimum rooted assembly space can be reduced to a grammar compression scheme demonstrates not only that the assembly process’ complexity is dependent on such contingencies as which assembly spaces or which distinct paths were taken in past historical stages of building the object, but also that this recursivity and dependence on past trajectories are subsumed into the concept of compressibility. However, unlike algorithmic approaches, assembly theory does not incorporate any of the elements of causality such as perturbation and counterfactual analysis, as has been done in the context of algorithmic complexity [30], a generalisation that when compared to AT reflects the concealed simplistic nature of AT.